# Gradient-based Local Next-best-view Planning for Improved Perception of Targeted Plant Nodes

### v1

@@ Abstract
Robots are increasingly used in tomato greenhouses to automate labour-intensive tasks such as selective harvesting and de-leafing. To perform these tasks, robots must be able to accurately and efficiently perceive the plant nodes that need to be cut, despite the high levels of occlusion from other plant parts. We formulate this problem as a local next-best-view (NBV) planning task where the robot has to plan an efficient set of camera viewpoints to overcome occlusion and improve the quality of perception. Our formulation focuses on quickly improving the perception accuracy of a single target node to maximise its chances of being cut. Previous methods of NBV planning mostly focused on global view planning and used random sampling of candidate viewpoints for exploration, which could suffer from high computational costs, ineffective view selection due to poor candidates, or non-smooth trajectories due to inefficient sampling. We propose a gradient-based NBV planner using differential ray sampling, which directly estimates the local gradient direction for viewpoint planning to overcome occlusion and improve perception. Through simulation experiments, we showed that our planner can handle occlusions and improve the 3D reconstruction and position estimation of nodes equally well as a sampling-based NBV planner, while taking ten times less computation and generating $28\%$ more efficient trajectories.
@@ I Introduction
Robots are increasingly used in agro-food environments, such as tomato greenhouses, to meet the growing demand for food and to compensate for the growing labour shortage [1, 2]. Using robots, we can help automate labour-intensive tasks like harvesting and de-leafing in a greenhouse. To perform these tasks, robots must accurately and efficiently perceive the plant nodes, i.e., the plant parts that connect the leaves and fruits to the main stem. Detecting these nodes and localising their 3D position is essential for planning a cutting and grasping action for harvesting and de-leafing. However, this is extremely challenging due to the high levels of occlusion in tomato plants, as the nodes are often hidden from the robot’s view either by leaves or other plant parts [3, 4].
@@
Active vision is a promising approach to handling occlusion, in which the robot reasons on the information gathered so far and strategically moves the camera to better viewpoints to overcome occlusion and improve perception accuracy [5]. Most active-vision algorithms are designed for global view planning, i.e. they aim to perceive an entire scene or a large structure. However, there are several tasks in a greenhouse that would benefit from a more local view planning that aims to perceive a specific plant part in more detail, such as inspection of fruits or pose estimation of cutting points. In such tasks, the goal is to plan viewpoints around a single target object and efficiently explore its properties. The widely-used approaches to global planning randomly sample a set of candidates to explore potential viewpoints for improving perception. Such global sampling-based approaches have some drawbacks for local planning: (i) a large number of candidates need to be sampled to explore the viewing space sufficiently, (ii) if insufficient candidates are sampled, it is highly likely that the actual best viewpoint is never sampled and hence never visited, and (iii) the trajectory of viewpoints generated may not be smooth and could have abrupt transitions across the viewing space. These drawbacks are more severe when the target object is highly occluded or the candidates are sampled with a greater degree-of-freedom (DoF) as the search space grows exponentially. As a result, sampling-based planning could result in longer processing times, higher computation costs, and an inefficient viewpoint trajectory for local planning.
@@
In this paper, we aimed to address these drawbacks of sampling-based approaches for local viewpoint planning. We developed a gradient-based optimisation algorithm by using a differentiable utility function, which allowed us to move the camera such that the viewpoint utility was locally maximised. For the perception of a single target object, a local optima is desirable as it locally improves coverage and handles occlusion. We applied our method to the problem of 3D reconstruction and position estimation of plant nodes. Using a simulated environment with 3D mesh models of tomato plants, we evaluated the performance of our method and compared it against a sampling-based and a random method for local view planning. The main contributions of this paper are: (i) a novel gradient-based approach for local view planning for handling occlusions and improving perception of a target object, (ii) significant improvements in computational speed and efficiency of the generated viewpoint trajectories compared to sampling-based and random methods, and (iii) the evaluation and analysis of our method in simulation using tomato plant models with high levels of occlusion.
@@ II Related work
Active vision is a well-researched topic [6], especially for exploration or reconstruction of unknown scenes. Next-best-view (NBV) planning [7] is a greedy approach to active vision, which plans one-step ahead. NBV planning methods can be categorised based on their viewpoint exploration strategy.
Frontier-based methods identify the boundaries between known and unknown regions of the scene and plan viewpoints to explore these frontiers [8]. They are primarily formulated for exploration of large scenes and hence cannot be directly used for other perception tasks. Sampling-based methods generate a set of candidate viewpoints and select the one that maximises a utility function as the next-best viewpoint. Rapidly-exploring Random Tree (RRT) is popular method for sampling the candidates [9]. These methods allow for fast exploration of the scene and can be adapted to other tasks such as localisation [10] and object search [11]. Hybrid approaches have also been proposed, that combine frontier-based and sampling-based methods by finding global frontiers and locally sampling candidates around them [12, 13]. Most of these methods are designed for global planning and are not suitable for local NBV planning for targeted perception.
Gradient-based methods are an alternate approach, which use the gradient of some viewpoint utility function to guide the selection of the next-best viewpoint. They can be beneficial for local planning as they avoid the computational cost of sampling candidate viewpoints and can generate a smooth sequence of viewpoints. Lehnert et al. [14] computed the gradient based on the surface area of the target object and showed that servoing the camera along the gradient improves the target perception. However, they used a special multi-camera setup to compute the gradients, making their method difficult to apply with other robotic platforms. Also, as they relied only on the current sensor data and did not explicitly map the scene. Deng et al. [15] proposed a fuzzy logic filter to make their frontier-based utility function differentiable. This allowed them to run a gradient-based optimisation to refine a global plan. However, their approach did not include semantic information to focus on target objects. Compared to these works, we propose a novel gradient-based approach for local NBV planning to perceive single target objects better. Our approach also uses a scene representation to merge information from multiple viewpoints, including semantic information.
@@ III Problem formulation
The common formulation of NBV planning methods is to find the next-best camera viewpoint to explore a bounded 3D space $V^{T}\subset\mathbb{R}^{3}$, to determine the free $V^{T}_{\text{free}}\subseteq V^{T}$ and occupied $V^{T}_{\text{occ}}\subseteq V^{T}$ regions, starting from $V^{T}$ being unknown. Since we are interested in perceiving only a target object, we define a region of interest (ROI) $V^{T}_{\text{ROI}}\subset V^{T}$, which is the region expected to contain the target object and is assumed to be given. Furthermore, we consider the semantic class labels and confidence scores from an object detection module to distinguish the target object from other objects. So, we modify the formulation of our NBV planning problem to finding the next-best camera viewpoint to explore $V^{T}$, to determine the free region $V^{T}_{\text{free}}$ and the region that belongs to the target object $V^{T}_{\text{tar}}\subset V^{T}$ within the ROI, starting from $V^{T}_{\text{ROI}}$ being unknown.
@@ III-A1 Optimisation parameters
The optimisation parameters depended on the viewpoint. The viewpoint was defined using a camera position $p^{c}\in V^{C}\subset\mathbb{R}^{3}$, where $V^{C}\cap V^{T}=\emptyset$, and an expected target position $p^{t}\in V^{T}$. We constrained the viewpoint such that the camera was positioned at $p^{c}$ and oriented towards $p^{t}$. The roll of the camera was left uncontrolled, as changing it would not contribute much to the utility. So, $p^{c}$ and $p^{t}$ defined a 5 degrees-of-freedom (DoF) viewpoint and hence formed the optimisation parameters,
@@ III-A3 Optimisation problem
The problem was to determine the viewpoint parameters that would maximise the expected utility $f(\xi)$, that is,
@@
s.t. $p^{c}\in V^{C},\;p^{t}\in V^{T}$ (3)
where $f(\xi)$ was the viewpoint utility and $V^{C}$ and $V^{T}$ defined the spatial constraints on the optimisation parameters. We hypothesise that such a gradient-based approach can smoothly guide the camera to perceive the target object better.
@@ IV Gradient-based local NBV planner
Figure 1: The pipeline of our proposed gradient-based NBV method. The nodes are detected in the color image using Mask R-CNN. The resulting semantic image, together with depth image, is inserted into the voxel grid to merge information from multiple views. The utility of the current view is computed using differentiable ray sampling, which provides a gradient along which the camera is moved.
To apply the method to our greenhouse scenario, we chose the target object as nodes of tomato plants, with the aim of reconstructing their 3D surface and estimating their 3D position. The major steps of our approach are as follows: (i) detect the node in the current viewpoint and estimate its position (Sec. IV-A), (ii) merge current information about the node with previously acquired information using a 4D scene representation (Sec. IV-B), (iii) compute the viewpoint utility of the current viewpoint and estimate the gradient (Sec. IV-C), and (iv) move in the direction of the gradient to potentially maximise the viewpoint utility (Sec. IV-D). We also keep track of the nodes over multiple viewpoints using an extended Kalman filter (Sec. IV-E). The pipeline of the proposed method is illustrated in Fig. 1.
@@ IV-B2 Semantic class label and probability
The semantic class label $c_{s}(x)$ referred to the type of node that a voxel $x$ belonged to, i.e. leaf or fruit, and the semantic probability $p_{s}(x)\in[0,1]$ referred to the probability of voxel $x$ belonging to the class $c_{s}(x)$. We also defined a background class to refer to voxels that did not belong to a node. Hence, the class labels were defined as $c_{s}(x)\in\{-1=\text{background},0=\text{fruit node},1=\text{leaf node}\}$. A semantic probability of $p_{s}=0$ implied that the voxel did not belong to class $c_{s}(x)$, while $p_{s}=1$ implied that the voxel belonged to class $c_{s}(x)$. $c_{s}(x)$ was most uncertain when the semantic probability was $0.5$. All voxels, except the ones within the ROI, were initialised as background with a probability close to zero, since we wanted to focus only on the target object and ignore the rest of the plant.
@@ IV-C Viewpoint utility with differentiable ray sampling
The utility of a viewpoint $\xi=\{p^{c},q^{c}\}$ was determined by estimating its semantic information gain $I_{s}(\xi)$, i.e., the amount of semantic information expected to be gained if the camera was moved there.
@@ IV-C1 Differentiable ray sampling
where $d$ was the viewing direction and $t$ limited the ray between the near $t_{n}=0.10$ and far $t_{f}=0.75$ bounds. Please note that the viewing direction $d$ depends on the camera orientation $q^{c}$ and hence $r$ is differentiable with respect to $\xi$. A set of $N_{r}=128$ points were uniformly sampled along each ray. For each sampled point, the occupancy and semantic probabilities were read from the voxel grid $\mathcal{M}$ by identifying the closest voxels to the sampled point and performing a trilinear interpolation of the occupancy and semantic probabilities of those voxels. The sampled probabilities were differentiable with respect to $\xi$, as the interpolation operation was implemented in a differentiable way, similar to [25].
@@ IV-D Viewpoint planning with gradient ascent
Once the gradient was computed, the viewpoint was moved in the direction of the gradient so that $I_{s}(\xi)$ could be maximised locally. This gradient-ascent step was scaled by a step size $\alpha$, which was a hyperparameter that needed to be tuned.
@@ IV-E Tracking nodes over multiple viewpoints
Since multiple nodes are detected by Mask R-CNN across multiple views, we needed a tracking method to properly associate the detected nodes in the current view with nodes from previous views, so that the position estimate of the target node could be correctly updated. We achieved this using an extended Kalman filter (EKF) with the 3D positions and class labels of nodes as states. Assuming that there were $N$ distinct nodes that were previously detected, the EKF state at step $k$ was defined as $S_{k}=\{o_{k}^{1},o_{k}^{2},...,o_{k}^{N}\}$, where $o_{k}^{j}=\{p_{k}^{j},c_{k}^{j}\}$ denoted the node $j$ and consisted of its 3D position $p_{k}^{j}$ and semantic class label $c_{k}^{j}$. The 3D position was defined with a Gaussian mean and covariance matrix, $p_{k}^{j}=\{\mu_{k}^{j},\Sigma_{k}^{j}\}$.
@@ IV-E2 State update
The 3D position was updated through a regular EKF update step which tries to minimise the mean squared error between the measured and estimated positions. For the class labels, we used a majority-voting-based approach, where the label that was observed in most of the viewpoints was assigned to the node.
V Baseline methods
@@
V-1 Sampling-based semantic next-best-view planner
@@ VI-A Simulation setup
[…]
@@ VI-B3 Number of ray-tracing calls
It indicated the computational cost of the view planners since the ray-tracing operation was the most computationally expensive step.
@@ VI-B5 Recall of occluded node detection
It measured the number of nodes that were detected accurately out of the subset of nodes that were occluded and undetected at the starting view. A detection was considered accurate when the 3D position of the node was within $0.02$m from the ground-truth position and the detected class label matched with the true class label.
VI-B6 Standard deviation of 3D node position
It was the estimated uncertainty on the node positions from the EKF. A low standard deviation indicated more confidence in the position estimation.

### v2

@@ Abstract
Robots are increasingly used in tomato greenhouses to automate labour-intensive tasks such as selective harvesting and de-leafing. To perform these tasks, robots must be able to accurately and efficiently perceive the plant nodes that need to be cut, despite the high levels of occlusion from other plant parts. We formulate this problem as a local next-best-view (NBV) planning task where the robot has to plan an efficient set of camera viewpoints to overcome occlusion and improve the quality of perception. Our formulation focuses on quickly improving the perception accuracy of a single target node to maximise its chances of being cut. Previous methods of NBV planning mostly focused on global view planning and used random sampling of candidate viewpoints for exploration, which could suffer from high computational costs, ineffective view selection due to poor candidates, or non-smooth trajectories due to inefficient sampling. We propose a gradient-based NBV planner using differentiable ray sampling, which directly estimates the local gradient direction for viewpoint planning to overcome occlusion and improve perception. Through simulation experiments, we showed that our planner can handle occlusions and improve the 3D reconstruction and position estimation of nodes equally well as a sampling-based NBV planner, while taking ten times less computation and generating $28\%$ more efficient trajectories.
@@ I Introduction
Robots are increasingly used in agro-food environments, such as tomato greenhouses, to meet the growing demand for food and to compensate for the growing labour shortage [1, 2]. Using robots, we can help automate labour-intensive tasks like harvesting and de-leafing in a greenhouse. To perform these tasks, robots must accurately and efficiently perceive the plant nodes, i.e., the plant parts that connect the leaves and fruits to the main stem. Detecting these nodes and localising their 3D position is essential to perform cutting and grasping action for harvesting and de-leafing. However, this is extremely challenging due to the high levels of occlusion in tomato plants, as the nodes are often hidden from the robot’s view by leaves or other plant parts [3, 4].
@@
Active vision is a promising approach to handling occlusion, in which the robot reasons based on the information gathered so far and strategically moves the camera to better viewpoints to overcome occlusion and improve perception accuracy [5]. Most active-vision algorithms are designed for global view planning, i.e. they aim to perceive an entire scene or a large structure. However, there are several tasks in a greenhouse that would benefit from a more local view planning that aims to perceive a specific plant part in more detail, such as inspection of fruits or pose estimation of cutting points. In such tasks, the goal is to plan viewpoints around a single target object and efficiently explore its properties. The widely-used approaches to global planning randomly sample a set of candidates to explore potential viewpoints for improving perception. Such global sampling-based approaches have some drawbacks for local planning: (i) a large number of candidates need to be sampled to explore the viewing space sufficiently, (ii) if insufficient candidates are sampled, it is highly likely that the actual best viewpoint is never sampled and hence never visited, and (iii) the trajectory of viewpoints generated may not be smooth and could have abrupt transitions across the viewing space. These drawbacks are more severe when the target object is highly occluded or the candidates are sampled with a greater degree-of-freedom (DoF) as the search space grows exponentially. As a result, sampling-based planning could result in longer processing times, higher computation costs, and an inefficient viewpoint trajectory for local planning.
@@
In this paper, we aimed to address these drawbacks of sampling-based approaches for local viewpoint planning. We developed a gradient-based optimisation algorithm by using a differentiable utility function, which allowed us to move the camera such that the viewpoint utility was locally maximised. For the perception of a single target object, a local optima is desirable as it locally improves coverage and handles occlusion. We applied our method to the problem of 3D reconstruction and position estimation of plant nodes. Using a simulated environment with 3D mesh models of tomato plants, we evaluated the performance of our method and compared it against a sampling-based and a random method for local view planning. The main contributions of this paper are: (i) a novel gradient-based approach for local view planning to handle occlusions and improve perception of a target object, (ii) significant improvements in computational speed and efficiency of the generated viewpoint trajectories compared to sampling-based methods, and (iii) the evaluation and analysis of our method in simulation using tomato plant models with high levels of occlusion. Our code is available at https://github.com/akshaykburusa/gradientnbv.
@@ II Related work
Active vision is a well-researched topic [6], especially for exploration or reconstruction of unknown scenes. Next-best-view (NBV) planning [7] is a popular approach to active vision, which plans one-step ahead. NBV planning methods can be categorised based on their exploration strategy.
Frontier-based methods identify the boundaries between known and unknown regions of the scene and plan viewpoints to explore these frontiers [8]. They are primarily formulated for exploration of large scenes and cannot be directly used for other perception tasks. Sampling-based methods generate a set of candidate viewpoints and select the one that maximises a utility function. Rapidly-exploring Random Tree (RRT) is popular method for sampling candidates [9]. These methods allow for fast exploration of the scene and can be adapted to tasks such as localisation [10] or object search [11]. Hybrid approaches have also been proposed, that combine frontier and sampling-based methods to globally find frontiers and locally sample candidates around them [12, 13]. These methods are designed for global planning and are not suitable for local or targeted NBV planning.
Gradient-based methods are an alternative approach, which use the gradient of some viewpoint utility function to guide the selection of the next-best viewpoint. They can be beneficial for local planning as they avoid the computational cost of sampling candidates and can generate a smooth sequence of viewpoints. In [14], it was shown that computing the gradient based on the surface area of a target object and servoing the camera along the gradient improved the target perception. However, they used a special multi-camera setup to compute the gradients, making their method difficult to apply on other platforms. Also, they relied only on the current sensor data and did not explicitly map the scene. In [15], a fuzzy logic filter was used to make a frontier-based utility function differentiable. This allowed the refinement of a global plan using gradient-based optimisation. However, their approach did not include semantic information to focus on target objects. Compared to these works, we propose a novel gradient-based approach for local NBV planning to perceive single target objects better, using volumetric and semantic information merged over multiple viewpoints.
@@ III Problem formulation
The common formulation of NBV planning methods is to find the next-best camera viewpoint to explore a bounded 3D space $V^{T}\subset\mathbb{R}^{3}$, to determine the free $V^{T}_{\text{free}}\subseteq V^{T}$ and occupied $V^{T}_{\text{occ}}\subseteq V^{T}$ regions, starting from $V^{T}$ being unknown. Since we are interested in perceiving only a target object, we define a region of interest (ROI) $V^{T}_{\text{ROI}}\subset V^{T}$, which is the region expected to contain the target object. We assume that the location of ROI within $V^{T}$ is given. Furthermore, we consider the semantic class labels and confidence scores from an object detection module to distinguish the target object from other objects. So, we modify the formulation of our NBV planning problem to finding the next-best camera viewpoint to explore $V^{T}_{\text{ROI}}$, to determine the free region $V^{T}_{\text{free}}$ and the region that belongs to the target object $V^{T}_{\text{tar}}\subset V^{T}_{\text{ROI}}$, starting from $V^{T}_{\text{ROI}}$ being unknown.
@@ III-A1 Optimisation parameters
The optimisation parameters depended on the viewpoint. The viewpoint was defined using a camera position $p^{c}\in V^{C}\subset\mathbb{R}^{3}$, where $V^{C}\cap V^{T}=\emptyset$, and an expected target position $p^{t}\in V^{T}_{\text{ROI}}$. We constrained the viewpoint such that the camera was positioned at $p^{c}$ and oriented towards $p^{t}$. The roll of the camera was left uncontrolled, as changing it would not contribute much to the utility. So, $p^{c}$ and $p^{t}$ defined a 5 degrees-of-freedom (DoF) viewpoint and hence formed the optimisation parameters,
@@ III-A3 Optimisation problem
The problem was to determine the viewpoint parameters that would maximise the expected viewpoint utility $f(\xi)$, that is,
@@
s.t. $p^{c}\in V^{C},\;p^{t}\in V^{T}_{\text{ROI}},$ (3)
where $V^{C}$ and $V^{T}_{\text{ROI}}$ defined the spatial constraints on the parameters. We hypothesise that such a gradient-based approach can smoothly guide the camera to perceive the target object better.
@@ IV Gradient-based local NBV planner
Figure 1: The pipeline of our proposed gradient-based NBV method. The nodes are detected in the color image using Mask R-CNN. The resulting segmented image, together with depth image, is inserted into the voxel grid to merge information from multiple views. The utility of the current view is computed using differentiable ray sampling, which provides a gradient along which the camera is moved.
To apply the method to our greenhouse scenario, we chose the target object as nodes of tomato plants, with the aim of reconstructing their 3D surface and estimating their 3D position. The major steps of our approach are as follows: (i) detect the node in the current viewpoint and estimate its position (Sec. IV-A), (ii) merge current information about the node with previously acquired information using a 4D scene representation (Sec. IV-B), (iii) compute the viewpoint utility of the current viewpoint and estimate the gradient (Sec. IV-C), and (iv) move in the direction of the gradient to potentially maximise the viewpoint utility (Sec. IV-D). We also keep track of the nodes over multiple viewpoints using a Kalman filter (Sec. IV-E). The pipeline of the proposed method is illustrated in Fig. 1.
@@ IV-B2 Semantic class label and probability
The semantic class label $c_{s}(x)$ referred to the type of node that a voxel $x$ belonged to, i.e. leaf or fruit, and the semantic probability $p_{s}(x)\in[0,1]$ referred to the probability of voxel $x$ belonging to the class $c_{s}(x)$. The voxels that did not belong to any node were considered as background. Hence, the class labels were defined as $c_{s}(x)\in\{-1=\text{background},0=\text{fruit node},1=\text{leaf node}\}$. A semantic probability of $p_{s}=0$ implied that the voxel did not belong to class $c_{s}(x)$, while $p_{s}=1$ implied that the voxel belonged to class $c_{s}(x)$. $c_{s}(x)$ was most uncertain when the semantic probability was $0.5$. All voxels, except the ones within the ROI, were initialised as background with a probability close to zero, since we wanted to focus only on the target object and ignore the rest of the plant.
@@ IV-C Viewpoint utility with differentiable ray sampling
The utility of a viewpoint $\xi=\{p^{c},p^{t}\}$ was determined by estimating its semantic information gain $I_{s}(\xi)$, i.e., the amount of semantic information expected to be gained if the camera was moved there.
@@ IV-C1 Differentiable ray sampling
where $d$ was the viewing direction and $t$ limited the ray between the near $t_{n}=0.10$ and far $t_{f}=0.75$ bounds. Please note that the viewing direction $d$ depends on the camera orientation $q^{c}$ and hence $r$ is differentiable with respect to $\xi$. A set of $N_{r}=128$ points were uniformly sampled along each ray. For each sampled point, the occupancy and semantic probabilities were read from the voxel grid $\mathcal{M}$ by identifying the closest voxels to the sampled point and performing a trilinear interpolation of the probabilities of those voxels. The sampled probabilities were differentiable with respect to $\xi$, as the interpolation operation was implemented in a differentiable way, similar to [25].
@@ IV-D Viewpoint planning with gradient ascent
Once the gradient was computed, the viewpoint was moved in the direction of the gradient so that $I_{s}(\xi)$ could be maximised locally. This gradient-ascent step was scaled by a step size $\alpha$, which was a hyperparameter.
@@ IV-E Tracking nodes over multiple viewpoints
Since multiple nodes are detected by Mask R-CNN across multiple views, we needed a tracking method to properly associate the detected nodes in the current view with nodes from previous views, so that the position estimate of the target node could be correctly updated. We achieved this using a Kalman filter with the 3D positions and class labels of nodes as states. Assuming that there were $N$ distinct nodes that were previously detected, the Kalman filter state at step $k$ was defined as $S_{k}=\{o_{k}^{1},o_{k}^{2},...,o_{k}^{N}\}$, where $o_{k}^{j}=\{p_{k}^{j},c_{k}^{j}\}$ denoted the node $j$ and consisted of its 3D position $p_{k}^{j}$ and semantic class label $c_{k}^{j}$. The 3D position was defined with a Gaussian mean and covariance matrix, $p_{k}^{j}=\{\mu_{k}^{j},\Sigma_{k}^{j}\}$. For targeted perception, one node from $S_{k}$ was selected as the target object and its position was assigned to $p^{t}$.
@@ IV-E2 State update
The 3D position was updated through a regular Kalman filter update step which tries to minimise the mean squared error between the measured and estimated positions. For the class labels, we used a majority-voting-based approach, where the label that was observed in most of the viewpoints was assigned to the node.
V Methods for comparison
@@
V-1 Sampling-based semantic NBV planner
@@ VI-A Simulation setup
The robot consisted of a 6-DoF manipulator (ABB IRB 1200) with an RGB-D camera (Intel RealSense L515) attached to it. The images obtained from the RGB-D camera had a resolution of $960\times 540$ pixels. The resolution of the voxel grid was $0.002$m and its dimensions were $0.3\times 0.3\times 0.7$m3. This dimension can be varied and does not constraint the algorithm. The step size for GradientNBV was set to $0.065$, which roughly moved the camera the same distance as SamplingNBV for the initial steps. We used eight 3D mesh models of tomato plants of varying growth stages and structural complexity.
@@ VI-B3 Number of ray-tracing calls
It indicated the computational cost of the view planners since the ray-tracing operation was the most computationally expensive step. The time for one ray-tracing call was $0.06$ seconds.
@@ VI-B5 Recall of occluded node detection
It measured the percentage of nodes that were detected accurately among the subset of nodes that were occluded and undetected at the starting view. A node was considered as accurately detected when its position was within $0.02$m from the ground-truth and its detected class label matched the true class label.
VI-B6 Standard deviation $(\sigma$) of 3D node position
It was the estimated uncertainty on the node position from the Kalman filter. A low standard deviation indicated more confidence in the position estimation.